# 1]FAIR at Meta 2]Univ. Grenoble Alpes, Inria, CNRS, Grenoble INP, LJK, France

### v1

@@ Appendix B Additional Experiment Results
B.1 Performance Between AR and OneFlow During Pretraining
@@
B.2 Pretraining from Scratch vs LLama Init
@@
B.3 Sampling Steps on Caption Performance
@@ Appendix D Full Derivations
we would need to marginalize out the auxiliary process $Z_{t}$ and the data $X_{1}$. Havasi et al. (2025) showed this can be done by using a loss based on any Bregman divergence while summing up over all possible sequences $z$ such that $x=f_{\text{rm-blanks}}(z)$. Concretely, given a convex function $\phi$ that defines a Bregman divergence $D_{\phi}(a,b)=\phi(a)-\phi(b)-\langle a-b,\tfrac{\mathrm{d}}{\mathrm{d}b}\phi(b)\rangle$, we can use the loss
@@ D.1 Interleaved time schedule
In order to model image insertions, we would make a choice. We can either (i) fully denoise images at the time of insertion, or (ii) insert only noise and denoise later. We choose the latter approach, as this allows simultaneous generation across images and text, and provides the best parallelism as only a single model forward at each step is needed for both modalities. Without loss of generality, assume there is only a single image.
Generation starts by advancing the sequence time, denoted $t_{\text{text}}=0$. When the image is inserted, we associate the image with its own time $t_{\text{img}}$.
@@
This will sample an extended time for the image in the internal [-1, 2]. If $\tau_{\text{img}}<0$, then it has not yet been inserted, hence it is deleted from the sequence. Otherwise, it is clip,

### v2

@@ Appendix B Additional Experiment Results
B.1 VQA Average Performance Across Categories
Figure 17: VQA performance across categories in controlled experiments. The largest gaps appear in General and OCR/Chart tasks.The models are tied on Vision-Centric benchmarks, which primarily depend on the visual encoder rather than text generation.
B.2 Performance Between AR and OneFlow During Pretraining
@@
B.3 Pretraining from Scratch vs LLama Init
@@
B.4 Sampling Steps on Caption Performance
@@ Appendix D Full Derivations
we would need to marginalize out the auxiliary process $Z_{t}$ and the data $X_{1}$. Havasi et al. (2025) showed this can be done by using a loss based on any Bregman divergence—following the recipe of Holderrieth et al. (2024)—while summing up over all possible sequences $z$ such that $x=f_{\text{rm-blanks}}(z)$. Concretely, given a convex function $\phi$ that defines a Bregman divergence $D_{\phi}(a,b)=\phi(a)-\phi(b)-\langle a-b,\tfrac{\mathrm{d}}{\mathrm{d}b}\phi(b)\rangle$, we can use the loss
@@ D.1 Interleaved time schedule
In order to model image insertions, we would make a choice. We can either (i) fully denoise images at the time of insertion, or (ii) insert only noise and denoise later. We choose the latter approach, as this allows simultaneous generation across images and text, and provides the best parallelism as only a single model forward at each step is needed for both modalities.
Without loss of generality, assume there is at most just one image. Generation starts by advancing the sequence time, denoted $t_{\text{text}}=0$. When the image is inserted, we associate the image with its own time $t_{\text{img}}$.
@@
This will sample an extended time for the image in the internal [-1, 2]. If $\tau_{\text{img}}<0$, then it has not yet been inserted, hence it is deleted from the sequence. Otherwise, it is clipped,